I need help solving this equation from my algebra 2 class.
it is a 3x3 matrix set equal to x^2. The instructions I've been given say, "Solve the equation."
the problem is as follows:
0 0 -1
3 x 0
2 0 3
equal to x^2

Answer by mananth(16949)

0 0 -1
3 x 0
2 0 3
...
0(3x-0)+0(9-0)+-1(0-2x)=x^2
0+0+2x=x^2
2x=x^2
x^2-2x=0
x(x-2)=0
x= 0 OR x=2